# Low dimensional dynamics of a sparse balanced synaptic network of quadratic integrate-and-fire neurons

### v1

@@ Abstract
Kinetics of a balanced network of neurons with a sparse grid of synaptic links is well representable by the stochastic dynamics of a generic neuron subject to an effective shot noise. The rate of delta-pulses of the noise is determined self-consistently from the probability density of the neuron states. Importantly, the most sophisticated (but robust) collective regimes of the network do not allow for the diffusion approximation, which is routinely adopted for a shot noise in mathematical neuroscience. These regimes can be expected to be biologically relevant. For the kinetics equations of the complete mean field theory of a homogeneous inhibitory network of quadratic integrate-and-fire neurons, we introduce circular cumulants of the genuine phase variable and derive a rigorous two cumulant reduction for both time-independent conditions and modulation of the excitatory current. The low dimensional model is examined with numerical simulations and found to be accurate for time-independent states and dynamic response to a periodic modulation deep into the parameter domain where the diffusion approximation is not applicable. The accuracy of a low dimensional model indicates and explains a low embedding dimensionality of the macroscopic collective dynamics of the network. The reduced model can be instrumental for theoretical studies of inhibitory-excitatory balances neural networks.
@@ 1 Introduction
For balanced networks with sparse grid of synaptic links, rich and nontrivial collective dynamics with important biological implications ware recently reported Goldobin-etal-2024 ; Goldobin-etal-2025 beyond the applicability limits of the diffusive approximation. Neither “next-generation neural mass models” not their upgraded versions for Gaussian noise cannot be employed for this case. In this paper we derive low dimensional model reductions on the basis of circular cumulants specifically for a population of quadratic integrate-and-fire neurons (QIFs) and a shot noise.
@@
where $S(t)$ is a Poissonian train of $\delta$ spikes with rate $R(t)=K\nu(t)$ and $\nu(t)$ is the population firing rate. For a homogeneous population within the mean-field framework the population dynamics can be described in terms of the membrane potential probability distribution function $P(V,t)$, whose time evolution is given by the continuity equation:
@@ 2 Continuity equation for the probability density of genuine phase
Eq. (4) cab be finally written in terms of $\psi$ as
@@
where $\min(n,m)$ stands for the minimal of two values, the binomial coefficients $\left(n\atop m\right)=\frac{n!}{m!(n-m)!}$; for obtaining the latter line one should separately consider two cases: $1\leq n\leq m$ and $n>m$. For convenience, we introduce
@@ 3 Two circular cumulant reduction for genuine phase
where the first CC is identical to the Kuramoto order parameter and the second CC measures the deviation of the second circular moment $z_{2}$ from the Ott–Antonsen manifold $z_{n}=(z_{1})^{n}$ Ott-Antonsen-2008 ; Ott-Antonsen-2009 . With a two circular cumulant reduction, the characteristic function $1+z_{1}k+z_{2}\frac{k^{2}}{2!}+z_{3}\frac{k^{3}}{3!}+z_{4}\frac{k^{4}}{4!}+\dots=e^{\varkappa_{1}k+\varkappa_{2}\frac{k^{2}}{2}}\approx(1+\varkappa_{2}\frac{k^{2}}{2})e^{\varkappa_{1}k}$; therefore, $z_{m}=\varkappa_{1}^{m}+\frac{m(m-1)}{2}\varkappa_{2}\varkappa_{1}^{m-2}$ Tyulkina-etal-2018 ; Goldobin-etal-2018 ; Goldobin-Dolmatova-2019 . Hence,
@@ 5 Circular cumulant equations with rescaled time
Figure 1: (a): The diagram of the macroscopic regimes (GOs: global oscillations, AS: asynchronous dynamics). The boundary between the GO and AS regimes is the Hopf bifurcation curve plotted with the black solid line for the original complete mean field system (6) and (31) and with the red dashed line for the diffusion approximation (Section 7). The cyan line: $K_{\mathrm{D}}(i_{0}/g_{0}^{2})$, the diffusion approximation is accurate for in- degree $K$ 1–2 orders of magnitude larger than $K_{\mathrm{D}}$ (Section 7.1). (b): Circular cumulants for the “exact” time-independent solution of (6) and (31) at the Hopf bifurcation line for $i_{0}/g_{0}^{2}=1.78$ (diamonds), $0.178$ (squares), $0.0316$ (triangles), $0.0007$ (black/blue circles for the upper/lower branch).
Figure 2: The firing rate for time-independent network states is plotted for $i_{0}/g_{0}^{2}=0.01$ (a), $0.03$ (b), $0.06$ (c), and $0.4$ (d). Black squares: the “exact” solution of the infinite $C$ equation chain (6) with firing rate (31); red diamonds: 2CC reduction (35)–(37); blue circles: Ott–Antonsen Ansatz given by Eqs. (35)–(36) with $\varkappa_{2}=0$. The values of $i_{0}$ and $K$ for these plots are shown with magenta lines in Figure 1a.
In Figure 2, one can see that 2CC reduction (35)–(37) provides reasonable accuracy for the time-independent states (AS) for as low excitatory current as $i_{0}=0.01g_{0}^{2}$, and the accuracy rapidly becomes better as $i_{0}$ increases. Here, for the sake of completeness, we also report the results for Ott–Antonsen (OA) Ansatz Ott-Antonsen-2008 ; Ott-Antonsen-2009 , which is given by Eqs. (35)–(36) with $\varkappa_{2}$ set to zero. One can see that the OA model reduction becomes accurate only for much higher values of the excitatory current $i_{0}$. Moreover, it is known to be fundamentally unable to reproduce the noise-induced oscillations in neural circuits Volo-Torcini-2018 ; Volo-etal-2022 ; Bi-Segneri-Volo-Torcini-2020 ; Goldobin-Volo-Torcini-2021 ; Ratas-Pyragas-2019 . The applicability and performance of the 2CC model reduction for the case of time-dependent synaptic activity are examined in the next Section.
@@ 8 Conclusion
For time-independent solutions the 2CC reduction is accurate for as low excitatory currents as $i_{0}=0.01g_{0}^{2}$, which extends by two orders of magnitude farther than the applicability domain of the diffusion approximation (Figure 1a). For self-excited noise-induced oscillations the 2CC reduction is less accurate since it underestimates the diffusive suppression of collective modes and gives a lowered value of the Hopf bifurcation threshold as compared to the “exact” solution; for heterogeneous populations the threshold is more accurate Goldobin-Volo-Torcini-2021 ; Volo-etal-2022 , but this is beyond the scope of our paper. Further, we analysed the accuracy of 2CC simulations for dynamic regimes (Figures 3 and 4). Even for a strong modulation of synaptic current with a resonant frequency, where the system is most sensitive to the problem of the overestimation of the noise-induced self-excitation of oscillations, the 2CC model captures the dynamic response for $i_{0}=0.2g_{0}^{2}$ and its error rapidly decreases as $i_{0}/g_{0}^{2}$ increases. For off-resonance frequencies, the dynamic response error is much smaller. Summarizing, the 2CC model can be expected to be applicable for theoretical studies of self-organized global oscillations in sparse networks where both excitatory and inhibitory synaptic links are present and collective oscillations emerge due to the interplay of excitation and inhibition Vreeswijk-Sompolinsky-1996 ; Kadmon-Sompolinsky-2015 ; Atay-etal-2011 ; Segneri-Bi-Olmi-Torcini-2020 ; Bi-Volo-Torcini-2021 ; Kirillov-Zlobin-Klinshov-2023 ; Kirillov-Smelov-Klinshov-2024 .
@@
The Ott–Antonsen reduction of the 2CC model has been found to be accurate only slightly beyond the applicability domain of the diffusion approximation. Moreover, it is completely unable to represent the effect on noise-induced self-excitation of collective oscillations. Comparing the 2CC reduction for DA (46)–(47) with the 2CC reduction for shot noise (40)–(40), one can see that the diffusion terms are different not only for the 2CC models but also for their OA reductions obtained by setting $\varkappa_{2}=0$.
@@ Acknowledgements
The authors are thankful to A. Torcini for discussions and acknowledge financial support by the Ministry of Science and Higher Education of the Russian Federation (theme no. 124021600038-9).
@@
All authors equally contributed to the study conception and design, theoretical derivations, and data analysis. Coding and numerical simulations were conducted by D.S. Goldobin. The first draft of the manuscript was written by M.V. Ageeva, the draft was finalized by D.S. Goldobin. All authors read and approved the final manuscript.

### v2

@@ Abstract
Kinetics of a balanced network of neurons with a sparse grid of synaptic links is well representable by the stochastic dynamics of a generic neuron subject to an effective shot noise. The rate of delta-pulses of the noise is determined self-consistently from the probability density of the neuron states. Importantly, the most sophisticated (but robust) collective regimes of the network do not allow for the diffusion approximation, which is routinely adopted for a shot noise in mathematical neuroscience. These regimes can be expected to be biologically relevant. For the kinetics equations of the complete mean field theory of a homogeneous inhibitory network of quadratic integrate-and-fire neurons, we introduce circular cumulants of the genuine phase variable and derive a rigorous two cumulant reduction for both time-independent conditions and modulation of the excitatory current. The low dimensional model is examined with numerical simulations and found to be accurate for time-independent states and dynamic response to a periodic modulation deep into the parameter domain where the diffusion approximation is not applicable. The accuracy of a low dimensional model indicates and explains a low embedding dimensionality of the macroscopic collective dynamics of the network. The reduced model can be instrumental for theoretical studies of inhibitory-excitatory balanced neural networks.
@@ 1 Introduction
For balanced networks with sparse grid of synaptic links, rich and nontrivial collective dynamics with important biological implications were recently reported Goldobin-etal-2024 ; Goldobin-etal-2025 beyond the applicability limits of the diffusive approximation. Neither “next-generation neural mass models” nor their upgraded versions for Gaussian noise can be employed for this case. In this paper we derive low dimensional model reductions on the basis of circular cumulants specifically for a population of quadratic integrate-and-fire neurons (QIFs) and a shot noise.
@@
where $S(t)$ is a Poissonian train of $\delta$ spikes with rate $R(t)=K\nu(t)$ and $\nu(t)$ is the population firing rate. For a homogeneous population within the mean-field framework the population dynamics can be described in terms of the membrane potential probability density function $P(V,t)$, whose time evolution is given by the continuity equation:
@@ 2 Continuity equation for the probability density of genuine phase
Eq. (4) can be finally written in terms of $\psi$ as
@@
where $\min(n,m)$ stands for the minimal of two values, the binomial coefficients $\left(n\atop m\right)=\frac{n!}{m!(n-m)!}$; for obtaining the line for $m\geq 1$ one should separately consider two cases: $1\leq n\leq m$ and $n>m$. For convenience, we introduce
@@ 3 Two circular cumulant reduction for genuine phase
where the first CC is identical to the Kuramoto order parameter and the second CC measures the deviation of the second circular moment $z_{2}$ from the Ott–Antonsen manifold $z_{n}=(z_{1})^{n}$ Ott-Antonsen-2008 ; Ott-Antonsen-2009 . With a two circular cumulant reduction, the characteristic function $1+z_{1}k+z_{2}\frac{k^{2}}{2!}+z_{3}\frac{k^{3}}{3!}+z_{4}\frac{k^{4}}{4!}+\dots\approx e^{\varkappa_{1}k+\varkappa_{2}\frac{k^{2}}{2}}\approx(1+\varkappa_{2}\frac{k^{2}}{2})e^{\varkappa_{1}k}$; therefore, $z_{m}\approx\varkappa_{1}^{m}+\frac{m(m-1)}{2}\varkappa_{2}\varkappa_{1}^{m-2}$ Tyulkina-etal-2018 ; Goldobin-etal-2018 ; Goldobin-Dolmatova-2019 . Hence,
@@ 5 Circular cumulant equations with rescaled time
Figure 1: (a): The diagram of the macroscopic regimes (GOs: global oscillations, AS: asynchronous dynamics). The boundary between the GO and AS regimes is the Hopf bifurcation curve plotted with the black solid line for the original complete mean field system (6) and (31) and with the red dashed line for the diffusion approximation (Section 7). The cyan line: $K_{\mathrm{D}}(i_{0}/g_{0}^{2})$, the diffusion approximation is accurate for in- degree $K$ 1–2 orders of magnitude larger than $K_{\mathrm{D}}$ (Section 7.1). The gray dash-dotted line: the Hopf bifurcation curve for model reduction (35)–(37). (b): Circular cumulants for the “exact” time-independent solution of (6) and (31) at the Hopf bifurcation line for $i_{0}/g_{0}^{2}=1.78$ (diamonds), $0.178$ (squares), $0.0316$ (triangles), $0.0007$ (black/blue circles for the upper/lower branch). (c): The Lyapunov exponent $\lambda$ of the time-independent solution within the Ott–Antonsen Ansatz [given by Eqs. (35)–(36) with $\varkappa_{2}=0$] is presented with the shadowgraph of $2\pi\sqrt{i_{0}}K^{1/4}\lambda/g_{0}^{2}$.
Figure 2: The firing rate for time-independent network states is plotted for $i_{0}/g_{0}^{2}=0.01$ (a), $0.03$ (b), $0.06$ (c), and $0.4$ (d). Black squares: the “exact” solution of the infinite equation chain (6) with firing rate (31); red diamonds: 2CC reduction (35)–(37); blue circles: Ott–Antonsen Ansatz given by Eqs. (35)–(36) with $\varkappa_{2}=0$. The values of $i_{0}$ and $K$ for these plots are shown with magenta lines in Figure 1a.
In Figure 2, one can see that 2CC reduction (35)–(37) provides reasonable accuracy for the time-independent states (AS) for as low excitatory current as $i_{0}=0.01g_{0}^{2}$, and the accuracy rapidly becomes better as $i_{0}$ increases. Here, for the sake of completeness, we also report the results for Ott–Antonsen (OA) Ansatz Ott-Antonsen-2008 ; Ott-Antonsen-2009 , which is given by Eqs. (35)–(36) with $\varkappa_{2}$ set to zero. One can see that the OA model reduction becomes accurate only for much higher values of the excitatory current $i_{0}$. Moreover, it is known to be fundamentally unable to reproduce the noise-induced oscillations in neural circuits within the diffusion approximation Volo-Torcini-2018 ; Bi-Segneri-Volo-Torcini-2020 ; Volo-etal-2022 ; Goldobin-Volo-Torcini-2021 ; Ratas-Pyragas-2019 . For the OA reduction of the shot-noise model, the Lyapunov exponent $\lambda$ is also always negative as shown in Figure 1c, where we explicitly indicate the asymptotic law which can be derived for large $i_{0}/g_{0}^{2}$ or large $K$: $\lambda_{\infty}\approx-g_{0}^{2}/(2\pi\sqrt{i_{0}}K^{1/4})$. The applicability and performance of the 2CC model reduction for the case of time-dependent synaptic activity are examined in the next Section.
@@ 6 Time-dependent external excitatory current $I(t)$
Figure 5: For the driving-free system, the firing rate is plotted versus $K$ at $i_{0}/g_{0}^{2}=0.2$ (a) and $0.4$ (b) for the “exact” solution (black squares), 2CC reduction (red diamonds), and OA Ansatz (blue circles). In panel (c), the firing rate oscillations are plotted for $K=1000$, $i_{0}/g_{0}^{2}=0.2$ (dashed lines) and $0.4$ (solid); black: the “exact” solution, red: 2CC model.
@@
In Figure 5, for the driving-free system, the self-excitation of collective oscillations via a supercritical Hopf bifurcation can be witnessed from the dependence of the firing rate on $K$. In Figure 1a, with the gray dash-dotted curve presenting the Hopf bifurcation threshold for the 2CC model reduction, one can see that the oscillatory instability threshold of this approximation is lowered in the parameter domain where the periodically driven oscillations are adequately reproduced by the 2CC model (i.e., for $i_{0}/g_{0}^{2}\gtrsim 0.2$).
@@ 8 Conclusion
For time-independent solutions the 2CC reduction is accurate for as low excitatory currents as $i_{0}=0.01g_{0}^{2}$, which extends by two orders of magnitude farther than the applicability domain of the diffusion approximation (Figure 1a). For self-excited noise-induced oscillations the 2CC reduction is less accurate since it underestimates the diffusive suppression of collective modes and gives a lowered value of the Hopf bifurcation threshold as compared to the “exact” solution; for heterogeneous populations the threshold is more accurate Goldobin-Volo-Torcini-2021 ; Volo-etal-2022 , but this is beyond the scope of our paper. Further, we analysed the accuracy of 2CC simulations for dynamic regimes (Figures 3 and 4). Even for a strong modulation of synaptic current with a resonant frequency, where the system is most sensitive to the problem of the overestimation of the noise-induced self-excitation of oscillations, the 2CC model captures the dynamic response for $i_{0}=0.2g_{0}^{2}$ and its error rapidly decreases as $i_{0}/g_{0}^{2}$ increases. For off-resonance frequencies, the dynamic response error is much smaller. Summarizing, the 2CC model might be expected to be applicable for theoretical studies of self-organized global oscillations in sparse networks where both excitatory and inhibitory synaptic links are present and collective oscillations emerge due to the interplay of excitation and inhibition Vreeswijk-Sompolinsky-1996 ; Kadmon-Sompolinsky-2015 ; Atay-etal-2011 ; Segneri-Bi-Olmi-Torcini-2020 ; Bi-Volo-Torcini-2021 ; Kirillov-Zlobin-Klinshov-2023 ; Kirillov-Smelov-Klinshov-2024 .
@@
The Ott–Antonsen reduction of the 2CC model has been found to be accurate only slightly beyond the applicability domain of the diffusion approximation. Moreover, it is completely unable to represent the effect on noise-induced self-excitation of collective oscillations (the Lyapunov exponent $\lambda$ in Figure 1c is always negative). Comparing the 2CC reduction for DA (46)–(47) with the 2CC reduction for shot noise (40)–(41), one can see that the diffusion terms are different not only for the 2CC models but also for their OA reductions obtained by setting $\varkappa_{2}=0$.
@@ Acknowledgements
The authors are thankful to A. Torcini for discussions. The work was carried out as part of a major scientific project (Agreement No. 075-15-2024-535 by 23 April 2024).
@@
All authors equally contributed to the study conception and design, theoretical derivations, coding, numerical simulations and data analysis. The first draft of the manuscript was written by M.V. Ageeva, the draft was finalized by D.S. Goldobin. All authors read and approved the final manuscript.